--You have purchased a small multi-family building in Montclair, NJ for $2M. You expect that NOI in your first year of ownership will be $175,000, what is the initial cap rate?

--Your acquisition was financed with a 10-year loan form $1.5M offering a 5% interest rate and a 20-year amortization period. If your marginal income tax rate is 30% and payments are made on an annual basis, what is the value of the mortgage interest deduction you receive in your first year of ownership?

--Assuming no capital expenditures, TI or leasing commissions, what is the equity dividend rate for the first year of ownership expressed in percentage terms?
Answer #1

Initial Capitalisation Rate = (175,000 / 2,000,000)×100

= 8.75%

Value of Mortgage interest deduction = 1500000×5/100

= 75,000

Equity Capital = $ 2million - $ 1.5 million = $0.5million

Dividend for equity = (175,000 - 75,000) × 70%

=70,000

Equity Dividend Rate = (70,000÷500,000) × 100

= 14%

Note : Interest and tax will be deducted from Net Operating Income to arrive at Dividend for equity.

Note : It is assumed that Dividend Payout Ratio is 100%